# FedADMM-InSa: An Inexact and Self-Adaptive ADMM for Federated Learning

### v1

@@ A.1.2 Optimality Conditions
For the point $w^{k+1}=\left(u^{k+1},z^{k+1},\lambda^{k+1}\right)^{\top}$ generated by (A.1.2), utilizing the definitions of $e_{i}^{k}(u_{i})$ in (3.6) and $e^{k}(u)$ in (A.2), and combining the first-order optimality condition of (A.7c), we have:
@@
These inequalities will be used later in the proof.
@@ A.1.3 Convergence Proof
$\sigma_{i}\left\|\beta_{i}(z^{k}-z^{k-1})+\lambda_{i}^{k}-\lambda_{i}^{k-1}\right\|$ (A.28)
@@
$=$ $\sigma_{i}\sqrt{\beta_{i}}\left(\beta_{i}\left\|z^{k}-z^{k-1}\right\|^{2}+\frac{1}{\beta_{i}}\left\|\lambda_{i}^{k}-\lambda_{i}^{k-1}\right\|^{2}+2\left(z^{k}-z^{k-1}\right)^{\top}\left(\lambda_{i}^{k}-\lambda_{i}^{k-1}\right)\right)^{\frac{1}{2}}$
@@
$\left\|v_{i}^{k}-v_{i}^{k-1}\right\|_{H_{\beta_{i}}}^{2}=\left(1+\delta\right)\beta_{i}\left\|z^{k}-z^{k-1}\right\|^{2}+\frac{1}{\beta_{i}}\left\|\lambda_{i}^{k}-\lambda_{i}^{k-1}\right\|^{2}+2\left(z^{k}-z^{k-1}\right)^{\top}\left(\lambda_{i}^{k}-\lambda_{i}^{k-1}\right).$ (A.29)
Substituting (A.28) back into (A.27) and using (A.29), we have
@@
$\sum_{k=1}^{K}\left(u^{k+1}-u\right)^{\top}\nabla_{u}L_{\alpha,\beta}\left(u^{k+1},z^{k},\lambda^{k}\right)$ (A.32)
@@
$+\sum_{i=1}^{m}\alpha_{i}\frac{1}{2\mu_{i}}\frac{\sigma_{i}-\sigma_{i}^{K+1}}{1-\sigma_{i}}\left(\left\|e_{i}^{0}\left(u_{i}^{1}\right)\right\|+\sqrt{\beta_{i}}\left\|v_{i}^{0}-v_{i}^{1}\right\|_{H_{\beta_{i}}}\right)^{2}$
@@
$+\sum_{k=1}^{K-1}\sum_{i=1}^{m}\alpha_{i}\frac{1}{2\mu_{i}}\frac{\sigma_{i}\beta_{i}}{1-\sigma_{i}}\left\|v_{i}^{k}-v_{i}^{k+1}\right\|_{{H_{\beta_{i}}}}^{2}-\sum_{k=1}^{K}\sum_{i=1}^{m}\frac{\alpha_{i}}{2}\left\|v_{i}^{k}-v_{i}^{k+1}\right\|_{{H_{\beta_{i}}}}^{2}$
@@
$+\frac{1}{2}\left(\left\|v^{1}-v\right\|_{H_{1}}^{2}-\left\|v^{K+1}-v\right\|_{H_{1}}^{2}-\left\|v^{K}-v^{K+1}\right\|_{H_{1}}^{2}\right).$ (A.34)
@@
$\sum_{k=1}^{K}\sum_{i=1}^{m}\alpha_{i}\left(c_{i}-\frac{\mu_{i}}{2}\frac{\sigma_{i}}{1-\sigma_{i}}\right)\left\|u_{i}^{k+1}-u_{i}^{*}\right\|^{2}+\sum_{k=1}^{K-1}\sum_{i=1}^{m}\frac{\alpha_{i}}{2}\left(1-\frac{1}{\mu_{i}}\frac{\sigma_{i}\beta_{i}}{1-\sigma_{i}}\right)\left\|v_{i}^{k}-v_{i}^{k+1}\right\|_{H_{\beta_{i}}}^{2}$ (A.35)
$\leq$ $\sum_{i=1}^{m}\alpha_{i}\frac{1}{2\mu_{i}}\frac{\sigma_{i}}{1-\sigma_{i}}\left(\left\|e_{i}^{0}\left(u_{i}^{1}\right)\right\|+\sqrt{\beta_{i}}\left\|v_{i}^{0}-v_{i}^{1}\right\|_{H_{\beta_{i}}}\right)^{2}$
@@
Therefore, the inequality (A.35) implies that

### v3

@@ 5 Conclusions
Acknowledgments
This work has been funded by the Humboldt Research Fellowship for postdoctoral researchers, the Alexander von Humboldt-Professorship program, the European Union’s Horizon Europe MSCA project ModConFlex (grant number 101073558), the COST Action MAT-DYN-NET, the Transregio 154 Project of the DFG, grants PID2020-112617GB-C22 and TED2021-131390B-I00 of MINECO (Spain). Madrid Government - UAM Agreement for the Excellence of the University Research Staff in the context of the V PRICIT (Regional Programme of Research and Technological Innovation).
@@ A.1.2 Optimality Conditions
For the point $w^{k+1}=\left(u^{k+1},z^{k+1},\lambda^{k+1}\right)^{\top}$ generated by (A.7), utilizing the definitions of $e_{i}^{k}(u_{i})$ in (3.6) and $e^{k}(u)$ in (A.2), and combining the first-order optimality condition of (A.7c), we have:
@@
The above inequalities will be used later in the proof.
@@ A.1.3 Convergence Proof
$\sigma_{i}\left\|\beta_{i}(z^{k}-z^{k-1})+\lambda_{i}^{k}-\lambda_{i}^{k-1}\right\|$
@@
$=$ $\sigma_{i}\sqrt{\beta_{i}}\left(\beta_{i}\left\|z^{k}-z^{k-1}\right\|^{2}+\frac{1}{\beta_{i}}\left\|\lambda_{i}^{k}-\lambda_{i}^{k-1}\right\|^{2}+2\left(z^{k}-z^{k-1}\right)^{\top}\left(\lambda_{i}^{k}-\lambda_{i}^{k-1}\right)\right)^{\frac{1}{2}}$ (A.28)
@@
$\left\|v_{i}^{k}-v_{i}^{k-1}\right\|_{H_{\beta_{i}}}^{2}=$ $\left(1+\delta\right)\beta_{i}\left\|z^{k}-z^{k-1}\right\|^{2}+\frac{1}{\beta_{i}}\left\|\lambda_{i}^{k}-\lambda_{i}^{k-1}\right\|^{2}$ (A.29)
$+2\left(z^{k}-z^{k-1}\right)^{\top}\left(\lambda_{i}^{k}-\lambda_{i}^{k-1}\right).$
Substituting (A.1.3) back into (A.27) and using (A.29), we have
@@
$\sum_{k=1}^{K}\left(u^{k+1}-u\right)^{\top}\nabla_{u}L_{\alpha,\beta}\left(u^{k+1},z^{k},\lambda^{k}\right)$
@@
$+\sum_{i=1}^{m}\alpha_{i}\frac{1}{2\mu_{i}}\frac{\sigma_{i}-\sigma_{i}^{K+1}}{1-\sigma_{i}}\left(\left\|e_{i}^{0}\left(u_{i}^{1}\right)\right\|+\sqrt{\beta_{i}}\left\|v_{i}^{0}-v_{i}^{1}\right\|_{H_{\beta_{i}}}\right)^{2}$ (A.32)
@@
$+\sum_{k=1}^{K-1}\sum_{i=1}^{m}\alpha_{i}\frac{1}{2\mu_{i}}\frac{\sigma_{i}\beta_{i}}{1-\sigma_{i}}\left\|v_{i}^{k}-v_{i}^{k+1}\right\|_{{H_{\beta_{i}}}}^{2}-\sum_{k=1}^{K}\sum_{i=1}^{m}\frac{\alpha_{i}}{2}\left\|v_{i}^{k}-v_{i}^{k+1}\right\|_{{H_{\beta_{i}}}}^{2}$ (A.34)
@@
$+\frac{1}{2}\left(\left\|v^{1}-v\right\|_{H_{1}}^{2}-\left\|v^{K+1}-v\right\|_{H_{1}}^{2}-\left\|v^{K}-v^{K+1}\right\|_{H_{1}}^{2}\right).$
@@
$\sum_{k=1}^{K}\sum_{i=1}^{m}\alpha_{i}\left(c_{i}-\frac{\mu_{i}}{2}\frac{\sigma_{i}}{1-\sigma_{i}}\right)\left\|u_{i}^{k+1}-u_{i}^{*}\right\|^{2}+\sum_{k=1}^{K-1}\sum_{i=1}^{m}\frac{\alpha_{i}}{2}\left(1-\frac{1}{\mu_{i}}\frac{\sigma_{i}\beta_{i}}{1-\sigma_{i}}\right)\left\|v_{i}^{k}-v_{i}^{k+1}\right\|_{H_{\beta_{i}}}^{2}$
$\leq$ $\sum_{i=1}^{m}\alpha_{i}\frac{1}{2\mu_{i}}\frac{\sigma_{i}}{1-\sigma_{i}}\left(\left\|e_{i}^{0}\left(u_{i}^{1}\right)\right\|+\sqrt{\beta_{i}}\left\|v_{i}^{0}-v_{i}^{1}\right\|_{H_{\beta_{i}}}\right)^{2}$ (A.35)
@@
Therefore, the inequality (A.1.3) implies that